In a class of 74 students, the number of females is one fewer than twice the number of males. How many females and how many males are there in the class
Answer by Susmitha(21)

given total no of students =74
a+b=74
given a is one fewer than twice the no of b
a= 2b-1
2b-1+b=74
3b-1=74
3b=74+1
3b=75
b=75/3
b=25
-------
a=2b-1
a=(2*25)-1
a=50-1
a=49
-------